OpenStudy (anonymous):

Will give medal! If you ate 1/4 of a pizza and there is 4 whole pieces left how many pieces of pizza were in the original pizza?

OpenStudy (anonymous):

kind of tricky isn't it?

OpenStudy (anonymous):

I'm think it is 5 whole pieces.

OpenStudy (anonymous):

you ate \(\frac{1}{4}\) of the pizza, that leaves \(\frac{3}{4}\) of it left 

OpenStudy (anonymous):

i think the problem is there were not a whole number of slices in the pizza

OpenStudy (anonymous):

OK.

OpenStudy (anonymous):

either that or i am interpreting the problem wrong when it says "there are 4 whole slices left" i take it to mean there are 4 whole slices, and not partial slices

OpenStudy (anonymous):

OK so the whole pizza had 4 pieces?

OpenStudy (anonymous):

three quarters of the pizza is 4 slices, so each piece is \(\frac{3}{4}\div 4=\frac{3}{16}\) of he whole pie

OpenStudy (anonymous):

2/16 is the answer?

OpenStudy (anonymous):

therefore there are \(\frac{16}{3}=5\tfrac{1}{3}\) pieces to begin with, which doesn't really make sense

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

hmm maybe i am messing up give me a second

OpenStudy (anonymous):

ok lets do it this way: call the number of slices in the whole pizza \(x\) then we have \(\frac{3}{4}\) of the pizza left, and we know \[\frac{3}{4}x=4\] this also gives \[x=\frac{16}{3}\]

OpenStudy (anonymous):

ok fine, it had five and a third slices